OpenStudy (anonymous):

what is the csc of 90 degrees?

OpenStudy (anonymous):

its 1

OpenStudy (trisaba):

1

OpenStudy (zzr0ck3r):

\(\csc(90)= \dfrac{1}{\sin(90)}=\dfrac{1}{?}\)

OpenStudy (anonymous):

tan= undefined sec= undefined

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

cot= 0

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

what about 180